# LLMs as Narcissistic Evaluators: When Ego Inflates Evaluation Scores

### v3

@@ 4.2 Bias towards longer summaries
Another notable pattern in Figure 2 is the high scores for the BART-based generators, indicated by both BARTScore variants and different GPTScores. To further investigate this phenomenon, we calculate the average length of summaries generated by each generator for each of the datasets. Notably, BART models and the Cohere that have not been fine-tuned for summarization tend to produce the longest summaries on average. This is followed by the fine-tuned BART models on the CNN dataset. Conversely, models fine-tuned on XSUM generate the shortest summaries. [^8] Note that the XSUM dataset contains one-sentence summaries.
@@
4.3 Qualitative Analysis: Correlation of Bias with Human Evaluation
To further verify the same model bias, we repeat the experiments from § 4.1 on summarization benchmarks that are accompanied by human evaluations. While the number of summaries in these benchmarks is limited compared to those in § 4.1, we can use the human annotations to verify that the inflated scores are not correlated with human evaluations.
Figure 4 shows the evaluation results for the SummEval and RoSE benchmarks for the reference-free and reference-based setting respectively. It is advisable to utilize the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes Yuan et al. (2021). Overall, we observe a trend similar to that shown in Figure 2. For instance, the T5-base generator receives higher scores from T5-based evaluators. [^9] In SummEval, the T5 model is only ranked higher when evaluated with certain variants of the T5Score in the reference-based setting. Meanwhile, BART-based models receive higher scores from both BARTScore and GPTScore evaluators, instead of T5 evaluator.
For human correlations, we use the RoSE and SummEval benchmark for evaluating the correlation in the reference-based and reference-free setting respectively. Table 3 (in Appendix) presents the Spearman and Kendall correlation values of SummEval in the reference-free setting. Overall, we observe that none of the evaluators have a strong correlation with the human annotations on either of these benchmarks. Due to the limited size of the samples (i.e., 100 summaries from SummEval and 100 summaries from ROSE with human annotations, as described in §3.3) and the absence of many of our investigated generators in § 4.1, we cannot draw a conclusive conclusion from the correlation values. Nevertheless, these results demonstrate that none of these evaluators highly correlate with human annotations, and as observed in § 4.1, their inflated scores for their own underlying generator may contribute to this low correlation.
Evaluator RoSE - Reference-based
@@ 4.3 Qualitative Analysis: Correlation of Bias with Human Evaluation
Table 2: Spearman and Kendall correlations between reference-based evaluation scores and human annotations using annotations in RoSE. The highest coefficient in each human annotation setting is bolded.
@@ 5 Conclusions
Based on experiments, we make the following conclusions:
First, the popularity of generative evaluation metrics, such as BARTScore, is on the rise for evaluating the factual accuracy of generated content—a critical concern in modern generator models. However, our results reveal that this evaluation approach is susceptible to the same model bias, highlighting the need for more robust metrics to assess factual correctness reliably. Second, our analysis indicates that models fine-tuned on the XSUM dataset are not suitable for direct integration into evaluators due to their bias towards shorter summaries. The exception is their use for evaluating summaries aligned with XSUM-style content. Third, notably, similar to traditional evaluation metrics Sun et al. (2019), contemporary evaluation metrics might also lean towards favoring longer summaries. This bias should be considered when interpreting and applying these metrics. Finally, our study uncovers the presence of the same model bias across all assessed evaluators. Consequently, we recommend avoiding the use of the same underlying model as the generator for assessment. Although the limited human evaluations for our examined models prevent definitive conclusions on selecting the best generative evaluator, our research charts a promising direction for designing more resilient and unbiased evaluation metrics.
@@
We noted that our work has the following limitations. Firstly, our experiment has been focused on the summarization task. Expanding the evaluation to encompass a broader range of generation tasks would be highly beneficial. Secondly, conducting a larger-scale human evaluation would be advantageous, as our current experiments are constrained by the limited sample sizes from SummEval and ROSE. Finally, incorporating additional generation models and evaluators in future work would further enrich the experiment.
@@
This paper raises no ethical concerns. The data and additional resources used in this study are open-source and widely employed in existing works.
@@ Appendix A Evaluation Setting
Evaluator SummEval - Reference-free
Spearman Kendall
BART-Base 0.023 0.010
BART-Large 0.120 0.082
BART-Base-CNN 0.240 0.164
BART-Large-CNN 0.227 0.149
BART-Base-XSUM 0.081 0.055
BART-Large-XSUM 0.061 0.043
T5-Small 0.396 0.275
T5-Base 0.365 0.249
T5-Large 0.337 0.236
T5-Small-CNN 0.448 0.315
T5-Base-CNN 0.259 0.172
T5-Small-XSUM 0.235 0.167
T5-Large-XSUM 0.110 0.081
GPT2 0.153 0.107
GPT2-Medium 0.207 0.146
GPT2-Large 0.183 0.128
GPT3-Curie 0.326 0.228
FLANT5-Base 0.361 0.252
FLANT5-XL 0.433 0.305
Cohere-Command 0.328 0.227
Table 3: Spearman and Kendall correlations between reference-free evaluation scores and human annotations using annotations in SummEval. The highest coefficient in each human annotation setting is bolded.
@@ A.1 Generator
Further information about models utilized in our generation models including checkpoint names, prompt setting, etc., are presented in Table 4.
@@
Table 4: Name of checkpoints or model utilized in our generation setting with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
@@ A.2 Evaluator
Further information about models utilized in our evaluation models are presented in Table 5.
@@
Table 5: Name of checkpoints or model utilized in our evaluation study with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
@@ B.1 Reference-free Setting
Figure 6: Assessing Bias in the XSUM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.

### v4

@@ 4.1 Quantitative Comparisons: Assessing Self-Bias in LM-Evaluators Towards Their Own Output
Figure 6: Assessing Bias in the XSUM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
@@ 4.2 Bias towards Longer Summaries
Another notable pattern in Figure 2 is the high scores for the BART-based generators, indicated by both BARTScore variants and different GPTScores. To further investigate this phenomenon, we calculate the average length of summaries generated by each generator for each of the datasets. Notably, BART models and Cohere that have not been fine-tuned for summarization tend to produce the longest summaries on average. This is followed by the fine-tuned BART models on the CNN dataset. Conversely, T5-based models score the summaries generated by Cohere low, as they tend to favour shorter summaries. A similar preference for short summaries can also be observed for evaluators fine-tuned on XSUM, which one-sentence summaries.
@@
4.3 Qualitative Analysis: Correlation of Self-Bias with Human Evaluation
To further verify the evaluators’ self-bias, we repeat the experiments from § 4.1 on summarization benchmarks that are accompanied by human evaluations. While the number of summaries in these benchmarks is limited compared to those in § 4.1, we can use the human annotations to verify that the inflated scores are not correlated with human evaluations.
Figure 4 shows the evaluation results for the SummEval and RoSE benchmarks for the reference-free and reference-based setting, respectively. As mentioned, we use SummEval for the reference-free setting and RoSE for the reference-based setting with regard to the specific aspects of each of these benchmarks Yuan et al. (2021). Overall, we observe a trend similar to that shown in Figure 2. For instance, the T5-base generator receives higher scores from T5-based evaluators. [^7] In SummEval, the T5 model is only ranked higher when evaluated with certain variants of the T5Score in the reference-based setting. Meanwhile, BART-based models receive higher scores from both BARTScore and GPTScore evaluators, instead of T5 evaluator.
Table 3 presents the Spearman and Kendall correlation values of SummEval in the reference-free setting, whereas the Spearman and Kendall correlation values of RoSE in the reference-based setting are given in Table 2.
Overall, we observe that none of the evaluators have a strong correlation with the human annotations on either of these benchmarks. Due to the limited size of the samples (i.e., 100 summaries from SummEval and 100 summaries from ROSE with human annotations, as described in §3.3) and the absence of many of our investigated generators in § 4.1, we cannot draw a conclusive conclusion from the correlation values. Nevertheless, these results demonstrate that none of these evaluators highly correlate with human annotations, and as observed in § 4.1, their inflated scores for their own underlying generator may contribute to this low correlation.
RoSE - Reference-based
Evaluator ACU
@@ 4.3 Qualitative Analysis: Correlation of Self-Bias with Human Evaluation
Table 2: Spearman and Kendall correlations between reference-based evaluation scores and human annotations using annotations in RoSE. Results in bold indicate the strongest coefficient.
SummEval - Reference-free
Evaluator Coherence Consistency Fluency Relevance
Spearman Kendall Spearman Kendall Spearman Kendall Spearman Kendall
BART-Base -0.028 -0.021 0.107 0.078 -0.043 -0.037 0.105 0.074
BART-Large 0.052 0.040 0.180 0.137 0.053 0.037 0.180 0.128
BART-Base-CNN 0.193 0.138 0.228 0.171 0.190 0.145 0.069 0.050
BART-Large-CNN 0.171 0.119 0.255 0.192 0.156 0.119 0.157 0.111
BART-Base-XSUM 0.170 0.120 -0.103 -0.079 0.068 0.055 -0.174 -0.124
BART-Large-XSUM 0.055 0.040 0.060 0.046 -0.025 -0.022 0.080 0.056
T5-Small 0.208 0.146 0.547 0.419 0.501 0.398 0.415 0.295
T5-Base 0.173 0.119 0.533 0.409 0.488 0.381 0.367 0.260
T5-Large 0.185 0.132 0.477 0.364 0.445 0.345 0.387 0.281
T5-Small-CNN 0.315 0.222 0.462 0.356 0.401 0.314 0.299 0.214
T5-Base-CNN 0.192 0.135 0.253 0.190 0.189 0.150 0.148 0.106
T5-Small-XSUM 0.245 0.178 0.142 0.109 0.209 0.164 0.113 0.079
T5-Large-XSUM 0.213 0.152 -0.111 -0.085 0.018 0.012 -0.041 -0.029
GPT2 0.103 0.077 0.154 0.117 0.037 0.026 0.032 0.021
GPT2-Medium 0.123 0.091 0.234 0.179 0.117 0.086 0.066 0.047
GPT2-Large 0.119 0.089 0.184 0.140 0.107 0.080 0.024 0.017
GPT3-Curie 0.152 0.108 0.483 0.371 0.345 0.264 0.311 0.223
FLANT5-Base 0.220 0.154 0.448 0.345 0.295 0.228 0.229 0.159
FLANT5-XL 0.248 0.174 0.550 0.424 0.389 0.301 0.402 0.289
Cohere-Command 0.136 0.097 0.520 0.397 0.351 0.268 0.427 0.302
Table 3: Spearman and Kendall correlations between the reference-free evaluation scores and expert annotations provided in SummEval on four different aspects. The strongest correlation for each aspect is bolded.
@@ 5 Conclusions
Based on experiments, we make the following conclusions: First, the popularity of generative evaluation metrics, such as BARTScore, is on the rise for evaluating the factual accuracy of generated content—a critical concern in modern generator models. However, our results reveal that this evaluation approach is susceptible to the self-bias, highlighting the need for more robust metrics to assess factual correctness reliably. Second, our analysis indicates that models fine-tuned on the XSUM dataset are not suitable for direct integration into evaluators due to their bias towards shorter summaries. The exception is their use for evaluating summaries aligned with XSUM-style content. Third, notably, similar to traditional evaluation metrics Sun et al. (2019), contemporary evaluation metrics might also lean towards favoring longer summaries. This bias should be considered when interpreting and applying these metrics. Finally, our study uncovers the presence of the self-bias across all assessed evaluators. Consequently, we recommend avoiding the use of the same underlying model as the generator for assessment. Although the limited human evaluations for our examined models prevent definitive conclusions on selecting the best generative evaluator, our research charts a promising direction for designing more resilient and unbiased evaluation metrics.
@@
We note that our work has the following limitations. Firstly, our experiment has been focused on the summarization task. Expanding the evaluation to encompass a broader range of generation tasks would be highly beneficial. Secondly, conducting a larger-scale human evaluation would be advantageous, as our current experiments are constrained by the limited sample sizes from SummEval and RoSE. Finally, incorporating additional generation models and evaluators in future work would further enrich the experiment.
@@
This paper raises no ethical concerns. The data and supplementary materials used in this study are open-sourced and widely employed in existing works.
@@ A.1 Generator
Full details of the models (e.g. checkpoint, prompt setting) that we employed as generators are given in Table 4.
@@
Table 4: Checkpoints or model utilized in our generation setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
@@ A.2 Evaluator
Full details of the models that we employed as our evaluators are given in Table 5 (reference-free settings) and Table 6 (reference-based settings).
@@
Table 5: Checkpoints or model utilized in our evaluation study for the reference-free setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
Name of Evaluator Name of Checkpoint or Model Suffix Prefix
BART-Base facebook/bart-base in other words:
BART-Large facebook/bart-large in other words:
BART-Base-CNN ainize/bart-base-cnn
BART-Large-CNN facebook/bart-large-cnn
BART-Base-XSUM morenolq/bart-base-xsum
BART-Large-XSUM facebook/bart-large-xsum
T5-Small t5-small Paraphrase:
T5-Base t5-base Paraphrase:
T5-Large t5-large Paraphrase:
T5-Small-CNN ubikpt/t5-small-finetuned-cnn
T5-Base-CNN flax-community/t5-base-cnn-dm
T5-Small-XSUM pki/t5-small-finetuned xsum
T5-Large-XSUM sysresearch101/t5-large-finetuned-xsum
GPT2 openai-community/gpt2 Paraphrase the sentence:
GPT2-Medium openai-community/gpt2-medium Paraphrase the sentence:
GPT2-Large openai-community/gpt2-large Paraphrase the sentence:
GPT3-Curie text-curie-001 Paraphrase the sentence:
FLANT5-Base google/flan-t5-base Paraphrase the sentence:
FLANT5-XL google/flan-t5-xl Paraphrase the sentence:
Cohere-Command api.cohere.ai/v1/generate Paraphrase the sentence:
Table 6: Checkpoints and models utilised in our evaluation study for the reference-based setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.